Calculate the price per share of common equity as of 12/31/2018.

 

 

 

 

 

 

Millions except price per share

Actual

 

 

 

 

 

2018/12/31

 

 

 

 

Value of operations (V0)

 

 

 

 

 

+ Value of short-term investments

 

 

 

 

 

Total value of company

 

 

 

 

 

Total value of all debt

 

 

 

 

 

Value of preferred stock

 

 

 

 

 

Value of common equity

 

 

 

 

 

Number of shares outstanding

 

 

 

 

 

Price per share

 

 

 

 

 

Income Statement for the Year Ending December 31 (Millions of Dollars) 2018 S Net Sales 800.0 Costs (except depreciation) Depreciation S 576.0 S 60.0 Total operating costs S 636.0 S Earning before int. & tax 164.0 S Less interest 32.0 Earning before taxes Taxes (40%) 132.0 52.8 S Net income before pref. div. 79.2 Preferred div. 1.4 Net income avail. for com. div. 77.9 S Common dividends 31.1 Addition to retained earnings 46.7 Number of shares (in millions) Dividends per share 10 3.11 Balance Sheets for December 31 (Millions of Dollars) Liabilities and Equity Accounts Payable Notes payable Assets 2018 2018 Cash 8.0 16.0 Short-term investments 20.0 40.0 80.0 Accounts receivable Accruals 40.0 Inventories 160.0 Total current liabilities $ 96.0 S Total current assets 268.0 Long-term bonds 300.0 Net plant and equipment 15.0 600.0 Preferred stock (Par plus PIC) S Total Assets 868.0 257.0 Retained earnings 200.0 S Common equity 457.0 Total liabilities and equit $ 868.0
Projected ratios and selected information for the current and projected years are shown below. Inputs Actual Projected Projected 2018/12/31 Projected Projected 12/31/21 12/31/19 12/31/20 12/31/22 Sales Growth Rate 15% 10% 6% 6% Costs/Sales 72% 72% 72% 72% 72% Depreciation/(Net PPE) 10% 10% 10% 10% 10% Cash/Sales 1% 1% 1% 1% 1% (Acct. Rec.)/Sales 10% 10% 10% 10% 10% Inventories/Sales 20% 20% 20% 20% 20% (Net PPE)/Sales (Acct. Pay.)/Sales 75% 75% 75% 75% 75% 2% 2% 2% 2% 2% Accruals/Sales 5% 5% 5% 5% 5% Tax rate 40% 40% 40% 40% 40% |Weighted average cost of capital (WACC) 10.5% 10.5% 10.5% 10.5% 10.5%
Answer #1

Price per share = $ 99.46

Please see the table below. Please be guided by the second column titled “Linkage” to understand the mathematics. You can also understand the mathematics making use of [+] / [-] sign appearing in the first column. The last row highlighted in yellow is your answer. Figures in parenthesis, if any, mean negative values. All financials are in $ mn except number of shares which is in mn and Price per share which is $ / shares.

Parameter Linkage Actual Projected Projected Projected Projected
Year 12/31/2018 12/31/2019 12/31/2020 12/31/2021 12/31/2022
Year number, n 1 2 3 4
y-o-y growth in sales gi 15% 10% 6% 6%
Net Sales A = Ai-1 x (1 + gi) 800         920.00       1,012.00       1,072.72       1,137.08
Costs B = 72% x A         662.40          728.64          772.36          818.70
Depreciation C = 10% x I (below)           69.00            75.90            80.45            85.28
Cash D = 1% x A                8.00             9.20            10.12            10.73            11.37
Acc. Receivables E = 10% x A              80.00           92.00          101.20          107.27          113.71
Inventory F = 20% x A            160.00         184.00          202.40          214.54          227.42
Accounts Payable G = 2% x A              16.00           18.40            20.24            21.45            22.74
Accruals H = 5% x A              40.00           46.00            50.60            53.64            56.85
Net PPE I = 75% x A            600.00         690.00          759.00          804.54          852.81
Working capital J = (D + E + F) - (G + H)            192.00         220.80          242.88          257.45          272.90
Increase in working capital K = Ji+1 - Ji           28.80            22.08            14.57            15.45
Capital Expenditure L = Ii + C - Ii-1         159.00          144.90          125.99          133.55
EBIT x (1 - Tax rate) M = (A - B - C) x (1 - 40%)         113.16          124.48          131.94          139.86
[+] Depreciation C           69.00            75.90            80.45            85.28
[-] Increase in working capital K           28.80            22.08            14.57            15.45
[-] Capital expenditure L         159.00          144.90          125.99          133.55
FCFF N = M + C - K - L            (5.64)            33.40            71.83            76.14
WACC R 10.5%
Terminal growth rate g 6%
Horizon value O = N2022 x (1 + g) / (R - g)       1,793.56
PV factor PVF = (1 + R)-n       0.90498        0.81898        0.74116        0.67073
PV of FCFF PV = PVF x FCFF            (5.10)            27.35            53.24            51.07
[+] PV of Horizon value PVHV = O x PVF2022       1,203.00
Value of operations (V0) V0 = PV + PVHV         1,329.56
[+] Value of short-term investments P              20.00
Total value of company Q = V0 + P         1,349.56
[-] Total value of all debt R (= Notes payables + long term bonds)            340.00
[-] Value of preferred stock S              15.00
Value of common equity T = Q - R - S            994.56
Number of shares outstanding (mn) U              10.00
Price per share ($ / share) V = T / U              99.46